# Can We Validate Counterfactual Estimations in the Presence of Unknown Network Interference?

### v1

@@ 9.5 Conditioning Technique
    $(\tilde{\rho}_{1N}{})^{2}$ $:=\mathrm{Var}\left[\Upsilon^{i}_{1}{}\Big{|}\vec{U}_{0}{}\right]=\frac{\sigma^{2}+\sigma^{2}_{0}}{N}\sum_{i=1}^{N}g_{0}\left(Y^{i}_{0}{},\vec{W}^{i},{\vec{X}}^{i}\right)^{2}.$ (38)
    Then, Assumption 9.4-(iii) implies that the value of $(\tilde{\rho}_{1N}{})^{2}$ is bounded independent from $N$, and
    $(\tilde{\rho}_{1}{})^{2}:=\lim_{N\rightarrow\infty}(\tilde{\rho}_{1N}{})^{2}.$ (39)
@@
    $\tilde{\nu}_{1N}^{i}{}$ $:=\frac{1}{N}\sum_{j=1}^{N}(\mu^{ij}+\mu_{0}^{ij})g_{0}\big{(}Y^{j}_{0}{},\vec{W}^{j},{\vec{X}}^{j}\big{)},\quad\quad\quad i\in[N].$ (41)
@@
    Above, we let $\bar{\mu}^{i}:=\mathbb{E}[{M}^{i}]$ and $\bar{\mu}^{i}_{0}:=\mathbb{E}[{M}_{0}^{i}]$, where ${M}^{i}\sim p_{\mu^{i}}$ and ${M}_{0}^{i}\sim p_{\mu^{i}_{0}}$. Additionally, $Y_{0}{}$, $\vec{W}$, and ${\vec{X}}$ represent the weak limits of the initial outcomes, treatment allocations, and covariates for the entire population as outlined in Assumption 9.4, respectively. In this context, $\bar{\mu}^{i}+\bar{\mu}^{i}_{0}$ determines the average interaction level of unit $i$ at time 00.
    Also, note that (42) yields the boundedness of $\tilde{\nu}_{1N}^{i}{}$ for all $i$ independent of $N$. Now, using Assmption 9.4 and the fact that $\psi\in{{\cal C}{\cal P}}(k)$, for $\kappa>0$, we have the following,
@@
    Then, applying Theorem 12.2 for the function $f(Y^{i}_{0}{},\vec{W}^{i},{\vec{X}}^{i},\tilde{\nu}_{1}^{i}{})=\mathbb{E}_{Z}\left[\psi\big{(}Y^{i}_{0}{},\tilde{\rho}_{1}{}Z,\vec{W}^{i},{\vec{X}}^{i},\tilde{\nu}_{1}^{i}{}\big{)}\right]$, we can write
@@
    The function $\widetilde{\psi}$ is within ${{\cal C}{\cal P}}(k)$ by Assumption 9.4. Then, applying (45) for the function $\widetilde{\psi}$, we obtain,
@@
    =$\;\stackrel{{\scriptstyle\mathclap{\mbox{\scriptsize a.s.}}}}{{=}}\;\mathbb{E}\Big{[}$ $\psi\big{(}Y_{0}{},\tilde{\nu}_{1}{}+\tilde{\rho}_{1}{}Z+H_{0}{},\tilde{\rho}_{1}{}Z,\bar{\rho}_{1}{}Z^{\prime},\vec{W},{\vec{X}},{M}^{j},{M}_{r}^{j}\big{)}\Big{]},$
@@
    =$\;\stackrel{{\scriptstyle\mathclap{\mbox{\scriptsize a.s.}}}}{{=}}\;\mathbb{E}\Big{[}$ $\psi\big{(}Y^{\mathcal{S}}_{0}{},\tilde{\nu}_{1}^{\mathcal{S}}{}+\tilde{\rho}_{1}{}Z+H^{\mathcal{S}}_{0}{},\tilde{\rho}_{1}{}Z,\bar{\rho}_{1}{}Z^{\prime},\vec{W}^{\mathcal{S}},{\vec{X}}^{\mathcal{S}},{M}^{\mathcal{S}},{M}_{r}^{\mathcal{S}}\big{)}\Big{]},$
@@
    $\lim_{N\rightarrow\infty}\frac{1}{N}\sum_{i=1}^{N}\big{(}\Upsilon^{i}_{1}{}\big{)}^{2}$ =$\;\stackrel{{\scriptstyle\mathclap{\mbox{\scriptsize a.s.}}}}{{=}}\;(\tilde{\rho}_{1}{})^{2}=\lim_{N\rightarrow\infty}\frac{\sigma^{2}+\sigma_{0}^{2}}{N}\sum_{i=1}^{N}\left(U^{i}_{0}{}\right)^{2}.$ (51)
@@
    $\lim_{N\rightarrow\infty}\frac{1}{N}\sum_{i=1}^{N}\big{(}\Upsilon^{i}_{1}{}-\vec{\text{B}}^{i\cdot}_{0}\vec{U}_{0}{}\big{)}^{2}$ =$\;\stackrel{{\scriptstyle\mathclap{\mbox{\scriptsize a.s.}}}}{{=}}\;(\bar{\rho}_{1}{})^{2}=\lim_{N\rightarrow\infty}\frac{\sigma^{2}}{N}\sum_{i=1}^{N}\left(U^{i}_{0}{}\right)^{2}.$ (52)
@@
    $\lim_{N\rightarrow\infty}\frac{1}{N}\sum_{i=1}^{N}\Upsilon^{i}_{1}{}\big{(}\Upsilon^{i}_{1}{}-\vec{\text{B}}^{i\cdot}_{0}\vec{U}_{0}{}\big{)}=\lim_{N\rightarrow\infty}\frac{1}{N}\sum_{i=1}^{N}\big{(}\vec{\text{A}}^{i\cdot}\vec{U}_{0}{}+\vec{\text{B}}^{i\cdot}_{0}\vec{U}_{0}{}\big{)}\big{(}\vec{\text{A}}^{i\cdot}\vec{U}_{0}{}\big{)}$ =$\;\stackrel{{\scriptstyle\mathclap{\mbox{\scriptsize a.s.}}}}{{=}}\;(\bar{\rho}_{1}{})^{2}=\lim_{N\rightarrow\infty}\frac{\sigma^{2}}{N}\sum_{n=1}^{N}\left(U^{n}_{0}{}\right)^{2}.$
@@
    For $t=1$, the matrix $\bm{Q}_{1}$ is equal to the vector $\vec{U}_{0}{}$ and $\bm{V}_{1}$ is equal to the vector $\vec{Y}_{1}{}$. By Assumption 9.4-(iii) and (51), we have
@@
=$\;\stackrel{{\scriptstyle\mathclap{\mbox{\scriptsize a.s.}}}}{{=}}\;\mathbb{E}\Big{[}$ $\psi\big{(}Y_{0}{},\tilde{\nu}_{1}{}+\tilde{\rho}_{1}{}Z_{1}+H_{0}{},\ldots,\tilde{\nu}_{t}{}+\tilde{\rho}_{t}{}Z_{t}+H_{t-1}{},\tilde{\rho}_{1}{}Z_{1},\ldots,\tilde{\rho}_{t}{}Z_{t},\bar{\rho}_{1}{}Z^{\prime}_{1},\ldots,\bar{\rho}_{t}{}Z^{\prime}_{t},\vec{W},{\vec{X}},{M}^{j},{M}_{r}^{j}\big{)}\Big{]},$
@@
=$\;\stackrel{{\scriptstyle\mathclap{\mbox{\scriptsize a.s.}}}}{{=}}\;\mathbb{E}\Big{[}$ $\psi\big{(}Y^{\mathcal{S}}_{0}{},\tilde{\nu}_{1}^{\mathcal{S}}{}+\tilde{\rho}_{1}{}Z_{1}+H^{\mathcal{S}}_{0}{},\ldots,\tilde{\nu}_{t}^{\mathcal{S}}{}+\tilde{\rho}_{t}{}Z_{t}+H^{\mathcal{S}}_{t-1}{},\tilde{\rho}_{1}{}Z_{1},\ldots,\tilde{\rho}_{t}{}Z_{t},\bar{\rho}_{1}{}Z^{\prime}_{1},\ldots,\bar{\rho}_{t}{}Z^{\prime}_{t},\vec{W}^{\mathcal{S}},{\vec{X}}^{\mathcal{S}},{M}^{\mathcal{S}},{M}_{r}^{\mathcal{S}}\big{)}\Big{]}.$
Also, for $10\leq s\neq r\leq t-$ 1, we have
$\lim_{N\rightarrow\infty}\frac{1}{N}\sum_{i=1}^{N}(\Upsilon^{i}_{r+1}{})^{2}$ =$\;\stackrel{{\scriptstyle\mathclap{\mbox{\scriptsize a.s.}}}}{{=}}\;(\tilde{\rho}_{r+1}{})^{2}\;\stackrel{{\scriptstyle\mathclap{\mbox{\scriptsize a.s.}}}}{{=}}\;\lim_{N\rightarrow\infty}\frac{\sigma^{2}+\sigma^{2}_{r}}{N}\sum_{i=1}^{N}(U^{i}_{r}{})^{2}$ (IH-3)
$\lim_{N\rightarrow\infty}\frac{1}{N}\sum_{i=1}^{N}\Upsilon^{i}_{r+1}{}\Upsilon^{i}_{s+1}{}$ =$\;\stackrel{{\scriptstyle\mathclap{\mbox{\scriptsize a.s.}}}}{{=}}\;\lim_{N\rightarrow\infty}\frac{\sigma^{2}}{N}\sum_{i=1}^{N}U^{i}_{r}{}U^{i}_{s}{},$ (IH-4)
$\lim_{N\rightarrow\infty}\frac{1}{N}\sum_{i=1}^{N}\big{(}\Upsilon^{i}_{r+1}{}-\vec{\text{B}}^{i\cdot}_{r}\vec{U}_{r}{}\big{)}^{2}$ =$\;\stackrel{{\scriptstyle\mathclap{\mbox{\scriptsize a.s.}}}}{{=}}\;(\bar{\rho}_{r+1}{})^{2}\;\stackrel{{\scriptstyle\mathclap{\mbox{\scriptsize a.s.}}}}{{=}}\;\lim_{N\rightarrow\infty}\frac{\sigma^{2}}{N}\sum_{i=1}^{N}\left(U^{i}_{r}{}\right)^{2},$ (IH-5)
$\lim_{N\rightarrow\infty}\frac{1}{N}\sum_{i=1}^{N}\big{(}\Upsilon^{i}_{r+1}{}-\vec{\text{B}}^{i\cdot}_{r}\vec{U}_{r}{}\big{)}\big{(}\Upsilon^{i}_{s+1}{}-\vec{\text{B}}^{i\cdot}_{s}\vec{U}_{s}{}\big{)}$ =$\;\stackrel{{\scriptstyle\mathclap{\mbox{\scriptsize a.s.}}}}{{=}}\;\lim_{N\rightarrow\infty}\frac{\sigma^{2}}{N}\sum_{i=1}^{N}U^{i}_{r}{}U^{i}_{s}{},$ (IH-6)
$\lim_{N\rightarrow\infty}\frac{1}{N}\sum_{i=1}^{N}\Upsilon^{i}_{r+1}{}\big{(}\Upsilon^{i}_{s+1}{}-\vec{\text{B}}^{i\cdot}_{s}\vec{U}_{s}{}\big{)}$ =$\;\stackrel{{\scriptstyle\mathclap{\mbox{\scriptsize a.s.}}}}{{=}}\;\lim_{N\rightarrow\infty}\frac{\sigma^{2}}{N}\sum_{i=1}^{N}U^{i}_{r}{}U^{i}_{s}{}.$ (IH-7)
@@
    We begin the second step by applying (IH-1) to the function $g_{s}\big{(}Y^{i}_{s}{},\vec{W}^{i},{\vec{X}}^{i}\big{)}g_{r}\big{(}Y^{i}_{r}{},\vec{W}^{i},{\vec{X}}^{i}\big{)}$, for $1\leq r,s\leq t$. Precisely, by Assumption 9.4-(iii) as well as (IH-1), we get
    $\lim_{N\rightarrow\infty}\frac{1}{N}\sum_{i=1}^{N}(U^{i}_{0}{})^{2}$ =$\;\stackrel{{\scriptstyle\mathclap{\mbox{\scriptsize a.s.}}}}{{=}}\;\mathbb{E}\left[g_{0}(Y_{0}{},\vec{W},{\vec{X}})^{2}\big{)}\right]=\frac{(\tilde{\rho}_{1}{})^{2}}{\sigma^{2}+\sigma_{0}^{2}}>0,$ (53)
    $\lim_{N\rightarrow\infty}\frac{1}{N}\sum_{i=1}^{N}U^{i}_{0}{}U^{i}_{s}{}$ =$\;\stackrel{{\scriptstyle\mathclap{\mbox{\scriptsize a.s.}}}}{{=}}\;\mathbb{E}\left[g_{0}(Y_{0}{},\vec{W},{\vec{X}})g_{s}\big{(}\tilde{\nu}_{s}{}+\tilde{\rho}_{s}{}Z_{s}+H_{s-1}{},\vec{W},{\vec{X}}\big{)}\right],$
    $\lim_{N\rightarrow\infty}\frac{1}{N}\sum_{i=1}^{N}U^{i}_{s}{}U^{i}_{r}{}$ =$\;\stackrel{{\scriptstyle\mathclap{\mbox{\scriptsize a.s.}}}}{{=}}\;\mathbb{E}\Big{[}g_{s}\big{(}\tilde{\nu}_{s}{}+\tilde{\rho}_{s}{}Z_{s}+H_{s-1}{},\vec{W},{\vec{X}}\big{)}g_{r}\big{(}\tilde{\nu}_{r}{}+\tilde{\rho}_{r}{}Z_{r}+H_{r-1}{},\vec{W},{\vec{X}}\big{)}\Big{]}.$
    Now, let $\vec{u}=(u_{1},\ldots,u_{t})^{\top}\in\mathbb{R}^{t}$ be a non-zero vector. By Assumption 9.4-(iii) and (53), we have
@@
    =$\;\stackrel{{\scriptstyle\mathclap{\mbox{\scriptsize a.s.}}}}{{=}}\;\mathbb{E}\left[\left(u_{1}g_{0}(Y_{0}{},\vec{W},{\vec{X}})+\sum_{s=1}^{t-1}u_{s+1}g_{s}\big{(}\tilde{\nu}_{s}{}+\tilde{\rho}_{s}{}Z_{s}+H_{s-1}{},\vec{W},{\vec{X}}\big{)}\right)^{2}\right]\geq 0.$
@@
    $\lim_{N\rightarrow\infty}\frac{1}{N}\sum_{i=1}^{N}\left(\vec{\widetilde{\text{A}}}^{i\cdot}\vec{U}^{\perp}_{t}{}\right)^{2}$ =$\;\stackrel{{\scriptstyle\mathclap{\mbox{\scriptsize a.s.}}}}{{=}}\;\lim_{N\rightarrow\infty}\frac{\sigma^{2}}{N}\sum_{i=1}^{N}\left(U^{\perp,i}_{t}{}\right)^{2},$ (57)
@@
    $\lim_{N\rightarrow\infty}\frac{1}{N}\sum_{i=1}^{N}\left(\vec{\text{B}}^{i\cdot}_{t}\vec{U}_{t}{}\right)^{2}$ =$\;\stackrel{{\scriptstyle\mathclap{\mbox{\scriptsize a.s.}}}}{{=}}\;\lim_{N\rightarrow\infty}\frac{\sigma_{t}^{2}}{N}\sum_{i=1}^{N}\left(U^{i}_{r}{}\right)^{2}.$ (58)
@@
    =$\;\stackrel{{\scriptstyle\mathclap{\mbox{\scriptsize a.s.}}}}{{=}}$ $\lim_{N\rightarrow\infty}\frac{\sigma^{2}}{N}\sum_{i=1}^{N}\sum_{0\leq s,r<t}\gamma^{t}_{s}\gamma^{t}_{r}U^{i}_{s}{}U^{i}_{r}{}$
@@
    $\lim_{N\rightarrow\infty}\frac{1}{N}\sum_{i=1}^{N}\big{(}\Upsilon^{i}_{t+1}{}\big{)}^{2}$ =$\;\stackrel{{\scriptstyle\mathclap{\mbox{\scriptsize a.s.}}}}{{=}}\;\lim_{N\rightarrow\infty}\frac{1}{N}\sum_{i=1}^{N}\left(\vec{\widetilde{\text{A}}}^{i\cdot}\vec{U}^{\perp}_{t}{}+\left[\bm{R}_{t}\vec{\gamma}_{t}\right]^{i}+\vec{\text{B}}^{i\cdot}_{t}\vec{U}_{t}{}\right)^{2}$ (60)
    =$\;\stackrel{{\scriptstyle\mathclap{\mbox{\scriptsize a.s.}}}}{{=}}\;\lim_{N\rightarrow\infty}\frac{\sigma^{2}}{N}\sum_{i=1}^{N}\left(U^{\perp,i}_{t}{}\right)^{2}+\lim_{N\rightarrow\infty}\frac{\sigma^{2}}{N}\sum_{i=1}^{N}\left(U^{\parallel,i}_{t}{}\right)^{2}+\lim_{N\rightarrow\infty}\frac{\sigma_{t}^{2}}{N}\sum_{i=1}^{N}\left(U^{i}_{t}{}\right)^{2}$

### v2

@@ 10.5 Conditioning Technique
    $(\tilde{\rho}_{1N}{})^{2}$ $=\mathrm{Var}\left[\Upsilon^{i}_{1}{}\Big|\vec{U}_{0}{}\right]=\frac{\sigma^{2}+\sigma^{2}_{0}}{N}\sum_{i=1}^{N}g_{0}\left(Y^{i}_{0}{},\vec{W}^{i},{\vec{X}}^{i}\right)^{2}.$ (43)
    Then, Assumption 8-(iii) implies that the value of $(\tilde{\rho}_{1N}{})^{2}$ is bounded independent from $N$, and
    $(\tilde{\rho}_{1}{})^{2}=\lim_{N\rightarrow\infty}(\tilde{\rho}_{1N}{})^{2}.$ (44)
@@
    $\tilde{\nu}_{1N}^{i}{}$ $=\frac{1}{N}\sum_{j=1}^{N}(\mu^{ij}+\mu_{0}^{ij})g_{0}\big(Y^{j}_{0}{},\vec{W}^{j},{\vec{X}}^{j}\big),\quad\quad\quad i\in[N].$ (46)
@@
    Above, we let $\bar{\mu}^{i}:=\mathbb{E}[{M}^{i}]$ and $\bar{\mu}^{i}_{0}:=\mathbb{E}[{M}_{0}^{i}]$, where ${M}^{i}\sim p_{\mu^{i}}$ and ${M}_{0}^{i}\sim p_{\mu^{i}_{0}}$; indeed, we used Assumption 10.4-(v) stating that ${M}^{i}$ and ${M}_{0}^{i}$ are independent of everything else. Additionally, $Y_{0}{}$, $\vec{W}$, and $\vec{X}$} represent the weak limits of the initial outcomes, treatment allocations, and covariates for the entire population as outlined in Assumption 10.4, respectively. In this context, $\bar{\mu}^{i}+\bar{\mu}^{i}_{0}$ determines the average interaction level of unit $i$ at time 0.
    Also, note that (47) yields the boundedness of $\tilde{\nu}_{1N}^{i}{}$ for all $i$ independent of $N$. Now, using Assmption 10.4 and the fact that $\psi\in{{\cal C}{\cal P}}(k)$, for $\kappa>0$, we have the following,
@@
    Then, applying Theorem 13.2 for the function $\text{f}(Y^{i}_{0}{},\vec{W}^{i},{\vec{X}}^{i},\tilde{\nu}_{1}^{i}{})=\mathbb{E}_{Z}\left[\psi\big(Y^{i}_{0}{},\tilde{\rho}_{1}{}Z,\vec{W}^{i},{\vec{X}}^{i},\tilde{\nu}_{1}^{i}{}\big)\right]$, we can write
@@
    The function $\widetilde{\psi}$ is within ${{\cal C}{\cal P}}(k)$ by Assumption 8. Then, applying (50) for the function $\widetilde{\psi}$, we obtain,
@@
    $\;\stackrel{{\scriptstyle\mathclap{\mbox{\scriptsize a.s.}}}}{{=}}\;\mathbb{E}\Big[$ $\psi\big(Y_{0}{},\tilde{\nu}_{1}{}+\tilde{\rho}_{1}{}Z+H_{0}{},\tilde{\rho}_{1}{}Z,\bar{\rho}_{1}{}Z^{\prime},\vec{W},{\vec{X}},{M}^{j},{M}_{r}^{j}\big)\Big],$
@@
    $\;\stackrel{{\scriptstyle\mathclap{\mbox{\scriptsize a.s.}}}}{{=}}\;\mathbb{E}\Big[$ $\psi\big(Y^{\mathcal{S}}_{0}{},\tilde{\nu}_{1}^{\mathcal{S}}{}+\tilde{\rho}_{1}{}Z+H^{\mathcal{S}}_{0}{},\tilde{\rho}_{1}{}Z,\bar{\rho}_{1}{}Z^{\prime},\vec{W}^{\mathcal{S}},{\vec{X}}^{\mathcal{S}},{M}^{\mathcal{S}},{M}_{r}^{\mathcal{S}}\big)\Big],$
@@
    $\lim_{N\rightarrow\infty}\frac{1}{N}\sum_{i=1}^{N}\big(\Upsilon^{i}_{1}{}\big)^{2}$ $\;\stackrel{{\scriptstyle\mathclap{\mbox{\scriptsize a.s.}}}}{{=}}\;(\tilde{\rho}_{1}{})^{2}=\lim_{N\rightarrow\infty}\frac{\sigma^{2}+\sigma_{0}^{2}}{N}\sum_{i=1}^{N}\left(U^{i}_{0}{}\right)^{2}.$ (56)
@@
    $\lim_{N\rightarrow\infty}\frac{1}{N}\sum_{i=1}^{N}\big(\Upsilon^{i}_{1}{}-\vec{\text{B}}^{i\cdot}_{0}\vec{U}_{0}{}\big)^{2}$ $\;\stackrel{{\scriptstyle\mathclap{\mbox{\scriptsize a.s.}}}}{{=}}\;(\bar{\rho}_{1}{})^{2}=\lim_{N\rightarrow\infty}\frac{\sigma^{2}}{N}\sum_{i=1}^{N}\left(U^{i}_{0}{}\right)^{2}.$ (57)
@@
    $\lim_{N\rightarrow\infty}\frac{1}{N}\sum_{i=1}^{N}\Upsilon^{i}_{1}{}\big(\Upsilon^{i}_{1}{}-\vec{\text{B}}^{i\cdot}_{0}\vec{U}_{0}{}\big)=\lim_{N\rightarrow\infty}\frac{1}{N}\sum_{i=1}^{N}\big(\vec{\text{A}}^{i\cdot}\vec{U}_{0}{}+\vec{\text{B}}^{i\cdot}_{0}\vec{U}_{0}{}\big)\big(\vec{\text{A}}^{i\cdot}\vec{U}_{0}{}\big)$ $\;\stackrel{{\scriptstyle\mathclap{\mbox{\scriptsize a.s.}}}}{{=}}\;(\bar{\rho}_{1}{})^{2}=\lim_{N\rightarrow\infty}\frac{\sigma^{2}}{N}\sum_{n=1}^{N}\left(U^{n}_{0}{}\right)^{2}.$
@@
    For $t=1$, the matrix $\bm{Q}_{1}$ is equal to the vector $\vec{U}_{0}{}$ and $\bm{V}_{1}$ is equal to the vector $\vec{Y}_{1}{}$. By Assumption 8-(iii) and (56), we have
@@
$\;\stackrel{{\scriptstyle\mathclap{\mbox{\scriptsize a.s.}}}}{{=}}\;\mathbb{E}\Big[$ $\psi\big(Y_{0}{},\tilde{\nu}_{1}{}+\tilde{\rho}_{1}{}Z_{1}+H_{0}{},\ldots,\tilde{\nu}_{t}{}+\tilde{\rho}_{t}{}Z_{t}+H_{t-1}{},\tilde{\rho}_{1}{}Z_{1},\ldots,\tilde{\rho}_{t}{}Z_{t},\bar{\rho}_{1}{}Z^{\prime}_{1},\ldots,\bar{\rho}_{t}{}Z^{\prime}_{t},\vec{W},{\vec{X}},{M}^{j},{M}_{r}^{j}\big)\Big],$
@@
$\;\stackrel{{\scriptstyle\mathclap{\mbox{\scriptsize a.s.}}}}{{=}}\;\mathbb{E}\Big[$ $\psi\big(Y^{\mathcal{S}}_{0}{},\tilde{\nu}_{1}^{\mathcal{S}}{}+\tilde{\rho}_{1}{}Z_{1}+H^{\mathcal{S}}_{0}{},\ldots,\tilde{\nu}_{t}^{\mathcal{S}}{}+\tilde{\rho}_{t}{}Z_{t}+H^{\mathcal{S}}_{t-1}{},\tilde{\rho}_{1}{}Z_{1},\ldots,\tilde{\rho}_{t}{}Z_{t},\bar{\rho}_{1}{}Z^{\prime}_{1},\ldots,\bar{\rho}_{t}{}Z^{\prime}_{t},\vec{W}^{\mathcal{S}},{\vec{X}}^{\mathcal{S}},{M}^{\mathcal{S}},{M}_{r}^{\mathcal{S}}\big)\Big].$
Also, for $0\leq s\neq r\leq t-1$, we have
$\lim_{N\rightarrow\infty}\frac{1}{N}\sum_{i=1}^{N}(\Upsilon^{i}_{r+1}{})^{2}$ $\;\stackrel{{\scriptstyle\mathclap{\mbox{\scriptsize a.s.}}}}{{=}}\;(\tilde{\rho}_{r+1}{})^{2}\;\stackrel{{\scriptstyle\mathclap{\mbox{\scriptsize a.s.}}}}{{=}}\;\lim_{N\rightarrow\infty}\frac{\sigma^{2}+\sigma^{2}_{r}}{N}\sum_{i=1}^{N}(U^{i}_{r}{})^{2}$ (IH-3)
$\lim_{N\rightarrow\infty}\frac{1}{N}\sum_{i=1}^{N}\Upsilon^{i}_{r+1}{}\Upsilon^{i}_{s+1}{}$ $\;\stackrel{{\scriptstyle\mathclap{\mbox{\scriptsize a.s.}}}}{{=}}\;\lim_{N\rightarrow\infty}\frac{\sigma^{2}}{N}\sum_{i=1}^{N}U^{i}_{r}{}U^{i}_{s}{},$ (IH-4)
$\lim_{N\rightarrow\infty}\frac{1}{N}\sum_{i=1}^{N}\big(\Upsilon^{i}_{r+1}{}-\vec{\text{B}}^{i\cdot}_{r}\vec{U}_{r}{}\big)^{2}$ $\;\stackrel{{\scriptstyle\mathclap{\mbox{\scriptsize a.s.}}}}{{=}}\;(\bar{\rho}_{r+1}{})^{2}\;\stackrel{{\scriptstyle\mathclap{\mbox{\scriptsize a.s.}}}}{{=}}\;\lim_{N\rightarrow\infty}\frac{\sigma^{2}}{N}\sum_{i=1}^{N}\left(U^{i}_{r}{}\right)^{2},$ (IH-5)
$\lim_{N\rightarrow\infty}\frac{1}{N}\sum_{i=1}^{N}\big(\Upsilon^{i}_{r+1}{}-\vec{\text{B}}^{i\cdot}_{r}\vec{U}_{r}{}\big)\big(\Upsilon^{i}_{s+1}{}-\vec{\text{B}}^{i\cdot}_{s}\vec{U}_{s}{}\big)$ $\;\stackrel{{\scriptstyle\mathclap{\mbox{\scriptsize a.s.}}}}{{=}}\;\lim_{N\rightarrow\infty}\frac{\sigma^{2}}{N}\sum_{i=1}^{N}U^{i}_{r}{}U^{i}_{s}{},$ (IH-6)
$\lim_{N\rightarrow\infty}\frac{1}{N}\sum_{i=1}^{N}\Upsilon^{i}_{r+1}{}\big(\Upsilon^{i}_{s+1}{}-\vec{\text{B}}^{i\cdot}_{s}\vec{U}_{s}{}\big)$ $\;\stackrel{{\scriptstyle\mathclap{\mbox{\scriptsize a.s.}}}}{{=}}\;\lim_{N\rightarrow\infty}\frac{\sigma^{2}}{N}\sum_{i=1}^{N}U^{i}_{r}{}U^{i}_{s}{}.$ (IH-7)
@@
    We begin the second step by applying (IH-1) to the function $g_{s}\big(Y^{i}_{s}{},\vec{W}^{i},{\vec{X}}^{i}\big)g_{r}\big(Y^{i}_{r}{},\vec{W}^{i},{\vec{X}}^{i}\big)$, for $1\leq r,s\leq t$. Precisely, by Assumption 8-(iii) as well as (IH-1), we get
    $\lim_{N\rightarrow\infty}\frac{1}{N}\sum_{i=1}^{N}(U^{i}_{0}{})^{2}$ $\;\stackrel{{\scriptstyle\mathclap{\mbox{\scriptsize a.s.}}}}{{=}}\;\mathbb{E}\left[g_{0}(Y_{0}{},\vec{W},{\vec{X}})^{2}\big)\right]=\frac{(\tilde{\rho}_{1}{})^{2}}{\sigma^{2}+\sigma_{0}^{2}}>0,$ (58)
    $\lim_{N\rightarrow\infty}\frac{1}{N}\sum_{i=1}^{N}U^{i}_{0}{}U^{i}_{s}{}$ $\;\stackrel{{\scriptstyle\mathclap{\mbox{\scriptsize a.s.}}}}{{=}}\;\mathbb{E}\left[g_{0}(Y_{0}{},\vec{W},{\vec{X}})g_{s}\big(\tilde{\nu}_{s}{}+\tilde{\rho}_{s}{}Z_{s}+H_{s-1}{},\vec{W},{\vec{X}}\big)\right],$
    $\lim_{N\rightarrow\infty}\frac{1}{N}\sum_{i=1}^{N}U^{i}_{s}{}U^{i}_{r}{}$ $\;\stackrel{{\scriptstyle\mathclap{\mbox{\scriptsize a.s.}}}}{{=}}\;\mathbb{E}\Big[g_{s}\big(\tilde{\nu}_{s}{}+\tilde{\rho}_{s}{}Z_{s}+H_{s-1}{},\vec{W},{\vec{X}}\big)g_{r}\big(\tilde{\nu}_{r}{}+\tilde{\rho}_{r}{}Z_{r}+H_{r-1}{},\vec{W},{\vec{X}}\big)\Big].$
    Now, let $\vec{u}=(u_{1},\ldots,u_{t})^{\top}\in\mathbb{R}^{t}$ be a non-zero vector. By Assumption 8-(iii) and (58), we have
@@
    $\;\stackrel{{\scriptstyle\mathclap{\mbox{\scriptsize a.s.}}}}{{=}}\;\mathbb{E}\left[\left(u_{1}g_{0}(Y_{0}{},\vec{W},{\vec{X}})+\sum_{s=1}^{t-1}u_{s+1}g_{s}\big(\tilde{\nu}_{s}{}+\tilde{\rho}_{s}{}Z_{s}+H_{s-1}{},\vec{W},{\vec{X}}\big)\right)^{2}\right]\geq 0.$
@@
    $\lim_{N\rightarrow\infty}\frac{1}{N}\sum_{i=1}^{N}\left(\vec{\widetilde{\text{A}}}^{i\cdot}\vec{U}^{\perp}_{t}{}\right)^{2}$ $\;\stackrel{{\scriptstyle\mathclap{\mbox{\scriptsize a.s.}}}}{{=}}\;\lim_{N\rightarrow\infty}\frac{\sigma^{2}}{N}\sum_{i=1}^{N}\left(U^{\perp,i}_{t}{}\right)^{2},$ (62)
@@
    $\lim_{N\rightarrow\infty}\frac{1}{N}\sum_{i=1}^{N}\left(\vec{\text{B}}^{i\cdot}_{t}\vec{U}_{t}{}\right)^{2}$ $\;\stackrel{{\scriptstyle\mathclap{\mbox{\scriptsize a.s.}}}}{{=}}\;\lim_{N\rightarrow\infty}\frac{\sigma_{t}^{2}}{N}\sum_{i=1}^{N}\left(U^{i}_{r}{}\right)^{2}.$ (63)
@@
    $\;\stackrel{{\scriptstyle\mathclap{\mbox{\scriptsize a.s.}}}}{{=}}$ $\lim_{N\rightarrow\infty}\frac{\sigma^{2}}{N}\sum_{i=1}^{N}\sum_{0\leq s,r<t}\gamma^{t}_{s}\gamma^{t}_{r}U^{i}_{s}{}U^{i}_{r}{}$
@@
    $\lim_{N\rightarrow\infty}\frac{1}{N}\sum_{i=1}^{N}\big(\Upsilon^{i}_{t+1}{}\big)^{2}$ $\;\stackrel{{\scriptstyle\mathclap{\mbox{\scriptsize a.s.}}}}{{=}}\;\lim_{N\rightarrow\infty}\frac{1}{N}\sum_{i=1}^{N}\left(\vec{\widetilde{\text{A}}}^{i\cdot}\vec{U}^{\perp}_{t}{}+\left[\bm{R}_{t}\vec{\gamma}_{t}\right]^{i}+\vec{\text{B}}^{i\cdot}_{t}\vec{U}_{t}{}\right)^{2}$ (65)
    $\;\stackrel{{\scriptstyle\mathclap{\mbox{\scriptsize a.s.}}}}{{=}}\;\lim_{N\rightarrow\infty}\frac{\sigma^{2}}{N}\sum_{i=1}^{N}\left(U^{\perp,i}_{t}{}\right)^{2}+\lim_{N\rightarrow\infty}\frac{\sigma^{2}}{N}\sum_{i=1}^{N}\left(U^{\parallel,i}_{t}{}\right)^{2}+\lim_{N\rightarrow\infty}\frac{\sigma_{t}^{2}}{N}\sum_{i=1}^{N}\left(U^{i}_{t}{}\right)^{2}$